OpenStudy (anonymous):

What is the factored form of the expression? s^4 – 16 A. (s - 2)^2(s + 2)^2 B. (s - 2)(s + 2) C. (s - i)(s + i)(s - 2)(s + 2) D. (s - 2i)(s + 2i)(s - 2)(s + 2)

OpenStudy (anonymous):

i think it is not (A)

Parth (parthkohli):

Think of \(s^4 - 16\) as \((s^2)^2 - (2^2)^2\). Now use difference of squares formula.

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

i think A

OpenStudy (anonymous):

Y

Parth (parthkohli):

\[ a^2 - b^2 = (a + b)(a - b) \implies (s^2)^2 - (2^2)^2 =(s^2 + 2^2)(s^2 - 2^2)\]No, @moha_10

OpenStudy (anonymous):

coz s is power to 4 so the corect answer is A as i think

OpenStudy (anonymous):

am i right ?@ParthKohli

Parth (parthkohli):

No @moha_10 We have \((s^2 + 2^2)(s^2 - 2^2)\) so far. Now we can factor \(s^2 - 2^2\) easily.

Parth (parthkohli):

\[a^2 -b^2 = (a + b)(a - b) \implies s^2 - 2^2 =\cdots \]

OpenStudy (anonymous):

you lost me with that

OpenStudy (anonymous):

okay parth by logic i can say A

OpenStudy (mathmate):

@moha_10 Follow @ParthKohli 's logic: a2−b2=(a+b)(a−b)⟹(s2)2−(22)2=(s2+22)(s2−22) (s^2-2^2) <=> (s+2)(s-2) But by the same token (s^2+2^2) can also be factorized into: (s^2+2^2) <=> (s+2i)(s-2i) where i^2=-1.

OpenStudy (anonymous):

\[s^4-16=\left(s^2+4\right)\left(s^2-4\right)=\left(s^2+4\right)(s-2) (s+2) \]